OpenStudy (anonymous):

which is the simplified form of 16q^5r^3/2q^4r^5

OpenStudy (anonymous):

Remember that \[\frac{ x^a }{ x^b } = x^{a - b}\]\[\dfrac{16q^5r^3}{2q^4r^5} = 8q^{5-4}r^{3-5}=8qr^{-2}= \frac{ 8q }{ r^2 }\]